### v1

@@ V.5.3 Pattern classifier
Beyond the distinctions in the waveform or time distribution, the planar distribution of signals within the top PMT arrays also presents notable differences between Pile-up DE and physical interactions. CENS and other physical signals are composed of multiple electrons generated from the same position above the liquid surface in the detector, while the DE signals are mostly generated around the previous muons’ tracks. These DEs are merged together into larger signals only due to their temporal proximity. Therefore, the distributions of photo-electrons from CENS signals and DE signals differ in the top PMT array of the detector. Real physical interactions appear more like a superposition of multiple single-electron signals from the same location, whereas the Pile-up DEs are formed by the superposition of single-electron signals originating from multiple different horizontal positions, thus appearing more dispersed.
After each S2 signal in the detector, the expected photons to be detected in each channel on the top PMT array are denoted as . is determined by the total area of S2, the relative LCE in that channel, and the PMT responses including detection efficiency. LCE of each PMT is derived with light propagation simulation in the RELICS TPC. It is assumed that the number of photo-electrons received by the PMT channel follows a Poisson distribution with as the mean. The detected photo-electrons are further derived considering the resolution of PMTs to single photo-election.
@@
We simulated 3.1 muon events, corresponding to 3.5 days of live time, and calculated the Pattern Likelihood and “Space-time Coefficient” . The two-dimensional distribution for Pile-up DE and physical interactions, within an S2 range of 120 to 240 PE, are shown in figure 13.
@@
While restricting the acceptance of Pile-up DE events lower than %, we scanned the parameter space of and , in order to retain as high signal acceptance as possible. The fitted values are of 9.10 , of 0.57 and of -65.78.
Figure 14 delineates the score distribution for Pile-up DEs and physical interactions. The final classifier identifies all events with scores exceeding 0 as real physical interactions. In the ROI of S2 ranging from 120 to 240 PE, the acceptance for physical interactions such as CENS is 67%. This cut reduced the rate of Pile-up DEs by 4 orders of magnitude in the same ROI, leaving the rate of Pile-up DE background below the expected CENS signal in the whole ROI.
@@ V.6 Summary
The background distribution from CRNs, cosmic muons, and detector materials are converted to S2 size from recoil energy using RelicsAPT which is described in Sec. III. All the background components mentioned above are summarized in figure 15. The CENS event rate is overlaid. The dominant background contribution comes from the Pile-up DEs, despite its large discrimination power to physical events as described in V.5. The NR background contribution from CRNs is an order of magnitude higher than muons, which is the second major contribution. The detector materials contribute a relatively low and flat component in ER. The event rate of all the background components in 32 kgyr exposure is summarized in Tab. 3. A signal-to-background ratio greater than 6 can be achieved in [120, 240] PEs ROI of CENS detection.
@@
| Pile-up DE | 1429.0 |
| CRNs | 470.4 |
| Muon-induced neutrons | 2.3 |
| ER | 232.0 |
| Total Background | 2133.7 |
| CENS | 13673.5 |
@@ VI.1 Weak mixing angle
The expected constraint on from RELICS is shown in figure 16. The great improvement from the red solid line to the blue dashed line indicates that the charge yield variation is a major contribution to the constraint uncertainty. The sensitivity is expected to be competitive and improve the Dresden-II’s result [61]. As illustrated in Fig. 17, RELICS will provide a measurement on with momentum transfer down to the MeV scale, which is a valuable complementary to the current measurements.
@@ VII Summary
The RELICS experiment is a (100) kg LXe detector with a fiducial mass of 32 kg that aims to observe the process of reactor CENS. The experiment is planned to be deployed at Sanmen Nuclear Power Plant (3 GW) in Zhejiang province, China, with a distance of 25 m, an average neutrino flux of 1013cm-2s-1.
In this work, a preliminary design of the RELICS experiment was presented. A dedicated MC simulation has been performed to study the background contributions. The dominant backgrounds come from the Pile-up DEs. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. An ionization-only analysis channel will be adopted for RELICS analysis to achieve a sufficiently low energy threshold for CENS detection. A signal-to-background ratio of greater than 6 can be achieved in [120, 240] PEs ROI of CENS detection. Its sensitivity to weak mixing angle and NSI are explored with 32 kgyr exposure.
@@
Acknowledgements. We acknowledge Sanmen Nuclear Power Plant for hosting RELICS. We thank Dr. Xun Chen for his support on the BambooMC package and Prof. Kaixuan Ni for fruitful discussions. RELICS is supported by a National Key R&D grant from the Ministry of Science and Technology of China (No. 2021YFA1601600) and a grant from Natural Science Foundation of China (No. 12275267). Yifei Zhao’s work on RELICS is supported by a grant for undergraduate research from Beijing Natural Science Foundation (No. QY23088).

### v2

@@ V.5.3 Pattern classifier
Beyond the distinctions in time distribution of the waveform, the planar distribution of signals within the top PMT arrays also presents notable differences between Pile-up DE and physical interactions. CENS and other physical signals are composed of multiple electrons generated from the same position below the liquid surface in the detector, while the DE signals are mostly generated around the previous muons’ tracks. These DEs are merged together into larger signals only due to their temporal proximity. Therefore, the distributions of photo-electrons from CENS signals and DE signals differ in the top PMT array of the detector. Real physical interactions appear more like a superposition of multiple single-electron signals from the same location, whereas the Pile-up DEs are formed by the superposition of single-electron signals originating from multiple different horizontal positions, thus appearing more dispersed.
After each S2 signal in the detector, the expected photons to be detected in each channel on the top PMT array are denoted as . is determined by the total area of S2, the relative LCE in that channel, and the PMT responses including detection efficiency. LCE of each PMT is derived with light propagation simulation in the RELICS TPC. It is assumed that the number of photo-electrons received by the PMT channel follows a Poisson distribution with as the mean. The detected photo-electrons are further derived considering the resolution of PMTs to a single photo-electron.
@@
We simulated the Pile-up DE events of the detector operating for 8500 hours, and calculated the “Pattern coefficient” and “Spacetime Correlation Coefficient” for each event. The two-dimensional distribution for Pile-up DE and CENS, within an S2 range of 150 to 160 PE, are shown in figure 13.
@@
While restricting the acceptance of Pile-up DE events lower than %, we scanned the parameter space of and , in order to retain as high signal acceptance as possible. The fitted values are of 5.6 , of 0.8 and of -37.42.
Figure 14 delineates the score distribution for Pile-up DEs and physical interactions. The final classifier identifies all events with scores exceeding 0 as real physical interactions. In the ROI of S2 ranging from 120 to 240 PE, the acceptance for physical interactions such as CENS is 46.5%. This cut reduced the rate of Pile-up DEs by 4 orders of magnitude in the same ROI, leaving the rate of Pile-up DE background below the expected CENS signal in the whole ROI.
@@ V.6 Summary
Considering the correlations between S2-width, waveform-classifier and pattern-classifier criteria, the total acceptance of these three cuts is estimated as shown in figure 15.
The distribution of all the background components is converted to S2 size from recoil energy using RelicsAPT, which is described in Sec. III, and summarized in figure 16. The CENS event rate is overlaid taking into account the signal acceptance. The dominant background contribution comes from the Pile-up DEs, despite its large discrimination power to physical events as described in Sec. V.5. The NR background contribution from CRNs is the second major contribution. The detector materials contribute a relatively low and flat component in ER. The background components’ event rate in 32 kgyr exposure is summarized in Tab. 3.
@@
| Pile-up DE | 1081.9 |
| CRNs | 229.4 |
| Muon-induced neutrons | 1.5 |
| ER | 5.7 |
| Total Background | 1318.4 |
| CENS | 4902.4 |
@@ VI.1 Weak mixing angle
The expected constraint on from RELICS is shown in figure 17. The great improvement from the red solid line to the blue dashed line indicates that the charge yield variation is a major contribution to the constraint uncertainty. The sensitivity is expected to be competitive and improve the Dresden-II’s result [63] measured with reactor neutrinos. As illustrated in figure 18, RELICS will provide a measurement on with momentum transfer down to the MeV scale, which is a valuable complementary to the current measurements.
@@ VII Summary
The RELICS experiment is a (100) kg LXe detector with a fiducial mass of 32 kg that aims to observe the process of reactor CENS. The experiment is planned to be deployed at the Sanmen Nuclear Power Plant (3 GW) in Zhejiang province, China, with a distance of 25 m, an average neutrino flux of 1013cm-2s-1.
In this work, a preliminary design of the RELICS experiment was presented. A dedicated MC simulation has been performed to study the background contributions. The dominant backgrounds come from the Pile-up DEs. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. An ionization-only analysis channel will be adopted for RELICS analysis to achieve a sufficiently low energy threshold for CENS detection. Its sensitivity to weak mixing angle and neutrino NSI are explored with 32 kgyr exposure.
@@
Acknowledgements. We acknowledge CNNC Sanmen Nuclear Power Company for hosting RELICS. We thank Dr. Xun Chen for supporting the BambooMC package and Prof. Kaixuan Ni for fruitful discussions. RELICS is supported by a National Key R&D grant from the Ministry of Science and Technology of China (No. 2021YFA1601600) and a grant from the Natural Science Foundation of China (No. 12275267). Yifei Zhao’s work on RELICS is supported by a grant for undergraduate research from Beijing Natural Science Foundation (No. QY23088).